OpenStudy (anonymous):

Find the indefinite integral: ∫ x^(2)e^(x) dx...

OpenStudy (anonymous):

=x^2∫ e^x -∫ (∫ e^x) d/dx x^2 =x^2 e^x - 2 ∫ e^x * x =x^2 e^x - 2 [ x∫ e^x - ∫∫e^x d/dx x] =x^2 e^x - 2 [ x e^x - ∫e^x] x^2 e^x - 2 x e^x + 2e^x +Constant